# Alternative ranking measures to predict international football results

### v2

@@ Abstract
Over the last few years, there has been a growing interest in the prediction and modelling of competitive sports outcomes, with particular emphasis placed on this area by the Bayesian statistics and machine learning communities. In this paper, we have carried out a comparative evaluation of statistical and machine learning models to assess their predictive performance for the 2022 FIFA World Cup and for the 2023 CAF Africa Cup of Nations by evaluating alternative summaries of past performances related to the involved teams. More specifically, we consider the Bayesian Bradley-Terry-Davidson model, which is a widely used statistical framework for ranking items based on paired comparisons that have been applied successfully in various domains, including football. The analysis was performed including in some canonical goal-based models both the Bradley-Terry-Davidson derived ranking and the widely recognized Coca-Cola FIFA ranking commonly adopted by football fans and amateurs.
@@ 1 Introduction
The application of statistical and machine learning models in forecasting international football competitions, such as FIFA World Cup or UEFA Champions League, has always attracted the interest of several analysts. From a statistical perspective, the outcome of a football match may be predicted using two different approaches. The result-based approach (Koning,, 2000; Carpita et al.,, 2019, among others), which directly predicts the match result, explicitly modelling the so-called three-way process – home win, draw, or away win – using a logistic or multinomial regression model. Alternatively, the goal-based approach models the goals scored and conceded in each match by modelling count variables, typically by using Poisson regressions, and then determines the exact match result by comparing these scores. It is worth noting that once a goal-based model has been estimated, it becomes possible to derive the three-way process by simply aggregating the estimated probabilities. However, for a more comprehensive comparison of goal-based and result-based statistical methods, see Egidi and Torelli, (2021)
In the goal-based approach, each game involves evaluating the goal counts for each team. Furthermore, the expected number of goals is determined by team attributes such as offensive and defensive abilities, and home advantage when applicable. Under some basic assumptions, the team-specific Poisson distributions are considered independent, resulting in a double-Poisson model (Maher,, 1982; Baio and Blangiardo,, 2010; Groll and Abedieh,, 2013; Egidi et al.,, 2018, among others). However, these Poisson-based approaches can be generalized in different ways to capture the dependence between scores. For instance, Dixon and Coles, (1997) extended the work of Maher, (1982) by allowing (a slightly negative) correlation between scores and incorporating a dependence parameter within their model to account for it. Furthermore, the bivariate Poisson model, designed to account for positive goal dependencies, was developed by Karlis and Ntzoufras, (2003) within a frequentist framework and by Ntzoufras, (2011) from a Bayesian perspective. A key limitation of previous models lies in their assumption of invariant team-specific parameters, implying that team performance remains constant over time based on their offensive and defensive abilities. However, it is recognised that team performance is inherently dynamic, fluctuating over years and possibly within seasons. Rue and Øyvind Salvesen, (2000) proposed a dynamic extension for the double Poisson model on continuous time, while Owen, (2011) proposed a discrete time random walk approach for both offensive and defensive parameters. In addition, Koopman and Lit, (2015) further extended the bivariate Poisson model into a state-space framework, allowing team abilities to vary according to a state vector.
A fundamentally different modelling approach has emerged due to the availability of large volumes of data, which has led to the development of a range of machine learning result-based techniques. These include artificial neural networks (ANNs), multivariate adaptive regression splines (MARS) (Friedman,, 1991) and ensemble learning methods such as random forests (Breiman,, 2001). Notably, the predictive performance of several random forests configurations have been examined and evaluated in the context of international football matches (Schauberger and Groll,, 2018; Groll et al.,, 2019, 2021, among others).
While the aforementioned methodologies provide robust frameworks for modelling football match outcomes, their predictive performance may be improved by incorporating additional historical information. Typically, one approach to potentially improve predictive accuracy is to integrate established rankings, such as the FIFA ranking, as additional model covariates. Notably, the algorithm underlying the FIFA ranking system has undergone a significant revision since 2018. The new algorithm takes into account not only the outcome of the single matches but also the strength level of teams before each match. For further details see Szczecinski and Roatis, (2022). The resulting algorithm offers benefits in terms of both simplicity and relative transparency. However, an alternative and particularly interesting methodology for obtaining a ranking, based on pairwise comparisons between teams, is the Bradley-Terry model (Bradley and Terry,, 1952). The model assigns a strength parameter to each team, and the odds of winning a match are determined by the ratio of these parameters. The estimated strengths can be employed to construct a rating system that reflects the ranking of teams based on the outcome of matches and the competitive interactions between teams.
In order to encompass a range of limitations, a number of extensions and generalisations to the original Bradley-Terry model have been developed over the years. For instance, Rao and Kupper, (1967) and Davidson, (1970) extended the model applicability to scenarios in which a draw is a possible outcome by including a novel parameter which affects the probability of a tie in a match. Springall, (1973) proposed a generalization of the Bradley-Terry model with team-dependent linear predictors. Davidson and Solomon, (1973) proposed a Bayesian version for the Bradley-Terry model, using a family of conjugate prior distributions to compute the posterior distribution of the log-strength parameters. Moreover, Leonard, (1977) suggested a more flexible Bayesian approach using non-conjugated multivariate Gaussian prior distributions for the log-strengths. Since these early works, there have been numerous contributions that have further extended the Bayesian framework for paired comparison models (Chen and Smith,, 1984; Caron and Doucet,, 2012; Whelan,, 2017; Osei and Davidov,, 2022; Wainer,, 2023, among others). The Bradley-Terry model can be further extended to account for situations where the order of comparisons can influence the outcome. A classic example is the home-field advantage in football, where the team playing at home may have a psychological or logistical advantage compared to the visiting team. Specifically, Beaver and Gokhale, (1975) and Davidson and Beaver, (1977) introduced additive and multiplicative order effects, respectively. Additionally, several extensions have been developed to model the dynamic variation of strengths over time. (Fahrmeir and Tutz,, 1994; Glickman,, 1999, 2001; Cattelan et al.,, 2012; Tian et al.,, 2023, among others).
@@
This paper aims to investigate the potential improvement in predictive performance achieved by some statistical goal-based methods and machine learning result-based algorithms when an appropriate ranking is used as an additional predictor. To asses it, we analyze the outcomes of the the most recent 2022 FIFA World Cup in Qatar and the 2023 Africa Cup in Ivory Coast. We considered a Bayesian Bradley-Terry-Davidson derived ranking system, using the posterior median of the log-strength parameters as novel predictor. Subsequently, we compare the predictive performances of this approach with those obtained using the well-established FIFA ranking to determine which of them provides more reliable forecasts. The rest of the paper is organized as follows. Section 2 delves into the theoretical framework, introducing the standard Bradley-Terry model and its Davidson extension for handling draws, followed by a discussion of the Bayesian approach. Furthermore, Section 3 describes the statistical goal-based methods and the machine learning result-based algorithms used in this study. In Section 4, we evaluate the application of these methodologies on the data from both the last World Cup and Africa Cup of Nations. Finally, Section 5 provides concluding remarks, outlining the limitations, advantages, and potential future research directions.
@@ 2 The Bradley-Terry Model
The Bradley-Terry model (Bradley and Terry,, 1952) is one of the most popular modelling techniques in a pairwise comparison context for ranking players or teams. The model assumes that each team $T_{i}$, with $i=1,\dots,N_{T}$, is characterized by a latent parameter, $\alpha_{i}>0$, representing its intrinsic strength. The outcome of any given comparison is modelled as an independent Bernoulli random variable, where the probability of each outcome is a function of the strengths of the teams involved. Specifically, for a match between team $T_{i}$ and team $T_{j}$, with $i\neq j=1,\dots,N_{T}$ , the probability that $T_{i}$ defeats $T_{j}$ in the n-th match, with $n=1,\dots,N$, is
@@
where $\alpha_{i}$ and $\alpha_{j}$ are the strength parameters of the teams involved in the match. These parameters are invariant to a multiplicative constant. Therefore, parameter identifiability is obtained by imposing a constraint such as $\sum_{k=1}^{N_{T}}\alpha_{k}=1$. Furthermore, the model in (1) is commonly reparameterized by the logarithm of the strength parameters
@@ 2.1 Dealing with draws
$p_{ij}^{D}$ $=\dfrac{\exp(\gamma+(\psi_{i}+\psi_{j})/2)}{\exp(\psi_{i})+\exp(\psi_{j})+\exp(\gamma+(\psi_{i}+\psi_{j})/2)}.$
It is worth to notice that, if the draw parameter $\gamma$ increases towards +$\infty$+ ∞ than the probability of a tie $p_{ij}^{D}$ approaches one. Conversely, if $\gamma$ decreases towards $\infty$- ∞ than $p_{ij}^{D}$ approaches to zero. Finally, if $\gamma$ is equal to zero then $p_{ij}^{D}$ and $p_{ij}^{W}$ depend solely on the strengths of the competing teams.
For the remainder of the paper, we will focus specifically on Davidson proposal for dealing with draws.
@@ 3 Statistical models and machine learning algorithms
This section describes the statistical models and machine learning algorithms employed to predict the outcomes of the considered competitions. Through a detailed analysis of goal-based models and result-based machine learning algorithms, this section aims to provide a comprehensive overview of the methodologies employed in the prediction of football matches, emphasising their statistical foundations and practical implementations in sports analytics.
@@ 3.1 Goal-based models
Goal-based models assume that the number of goals scored in a game by each team follows a discrete distribution, typically two independent Poisson or a bivariate-Poisson accounting for positive correlation. Thus, for each game, we need to consider the pair of counts $(X_{in},Y_{jn})$, for $i\neq j=1,\dots,N_{T}$ and $n=1,\dots,N$. The first count $X_{in}$ denotes the non-negative number of goals scored by the home team $T_{i}$ and the second count $Y_{jn}$ denotes the number of goals scored by the visiting team $T_{j}$, both in the n-th game. A simple double Poisson model is
@@
where $\lambda_{1n}$ and $\lambda_{2n}$ describe the expected number of goals for the home team and the away team, respectively. In particular, $\theta$ denotes a common baseline parameter, the parameters att and def represent the unknown attack and defense abilities for the home team $h_{n}$ and the away team $a_{n}$ in the n-th game. Furthermore, $\omega_{n}=(rank_{h_{n}}-rank_{a_{n}})$ captures the difference in rankings between the home and away teams in the n-th game. Finally, the parameter $\phi$ tries to correct for the ranking difference occurring between two competing teams. To ensure model identifiability, a sum-to-zero constraint (Baio and Blangiardo,, 2010) is imposed on the attack and defense parameters.
@@
Poisson goal-based models may suffer from an underestimation of the number of draws, represented by the outcomes in the diagonal of the probability table. To address this issue, Karlis and Ntzoufras, (2009) introduced a zero-inflated model for favouring the draw outcome. The diagonal-inflated bivariate Poisson model is defined as follows
@@
where $\mu_{\operatorname{att}}$ and $\mu_{\operatorname{def}}$ are the mean for the initial attack and defense abilities, and $\sigma_{\operatorname{att}}$ and $\sigma_{\operatorname{def}}$ are their corresponding variances. As the static models, the dynamic extension also imposes a sum-to-zero constraint on these random effects within each season for identifiability.
@@ 3.2 Result-based algorithms
Artificial neural networks (ANNs) represent a class of complex computational models inspired by the interconnected structure of neurons in the human brain. These models excel at processing information and learning from data through a layered architecture. Each layer comprises interconnected nodes that apply weights and biases to process inputs. The learning process involves adjusting these weights and biases to optimize the network’s performance. Specifically, ANNs proved successful in predicting football match outcomes by considering historical data that includes a wide range of information, such as team performance metrics, match results and even individual player statistics (Huang and Chang,, 2010; Hucaljuk and Rakipovic,, 2011; Danisik et al.,, 2018, among others).

### v3

@@ Abstract
Over the last few years, there has been a growing interest in the prediction and modelling of competitive sports outcomes, with particular emphasis placed on this area by the Bayesian statistics and machine learning communities. In this paper, we have carried out a comparative evaluation of statistical and machine learning models to assess their predictive performance for the 2022 FIFA World Cup and the 2023 CAF Africa Cup of Nations by evaluating alternative summaries of past performances related to the involved teams. More specifically, we consider the Bayesian Bradley-Terry-Davidson model, which is a widely used statistical framework for ranking items based on paired comparisons that have been applied successfully in various domains, including football. The analysis was performed including in some canonical goal-based models both the Bradley-Terry-Davidson derived ranking and the widely recognized Coca-Cola FIFA ranking commonly adopted by football fans and amateurs.
@@ 1 Introduction
The application of statistical and machine learning models in forecasting international football competitions, such as the FIFA World Cup or UEFA Champions League, has always attracted the interest of several analysts.
@@
From a statistical perspective, the outcome of a football match may be predicted using two different approaches. The result-based approach (Koning,, 2000; Carpita et al.,, 2019, among others), which directly predicts the match result, explicitly modelling the so-called three-way process – home win, draw, or away win – using a logistic or multinomial regression model. Alternatively, the goal-based approach models the goals scored and conceded in each match by modelling count variables, typically by using Poisson regressions, and then determines the exact match result by comparing these scores. It is worth noting that once a goal-based model has been estimated, it becomes possible to derive the three-way process by simply aggregating the estimated probabilities. However, for a more comprehensive comparison of goal-based and result-based statistical methods, see Egidi and Torelli, (2021).
@@
This paper investigates the potential improvement in predictive performance achieved by some statistical goal-based methods and machine learning result-based algorithms when an appropriate measure of the teams’ relative strength is used as an additional predictor. To assess it, we analyze the outcomes of the 2022 FIFA World Cup in Qatar and the 2023 Africa Cup in Ivory Coast. We considered a Bayesian Bradley-Terry-Davidson derived ranking system, using the posterior median of the log-strength parameters as a novel predictor. Subsequently, we compare the predictive performances of this approach with those obtained using the well-established FIFA ranking – in terms of FIFA ranking points – to determine which provides more reliable forecasts.
In the goal-based approach, each game involves evaluating the goal counts for each team. Furthermore, the expected number of goals is determined by team attributes such as offensive and defensive abilities, and home advantage when applicable. Under some basic assumptions, the team-specific Poisson distributions are considered independent, resulting in a double Poisson model (Maher,, 1982; Baio and Blangiardo,, 2010; Groll and Abedieh,, 2013; Egidi et al.,, 2018, among others). However, these Poisson-based approaches can be generalized in different ways to capture the dependence between scores. For instance, Dixon and Coles, (1997) extended the work of Maher, (1982) by allowing (a slightly negative) correlation between scores and incorporating a dependence parameter within their model to account for it. Furthermore, the bivariate Poisson model, designed to account for positive goal dependencies, was developed by Karlis and Ntzoufras, (2003) within a frequentist framework and by Ntzoufras, (2011) from a Bayesian perspective. A key limitation of previous models lies in their assumption of invariant team-specific parameters, implying that team performance remains constant over time based on their offensive and defensive abilities. However, it is recognised that team performance is inherently dynamic, fluctuating over years and possibly within seasons. Rue and Øyvind Salvesen, (2000) proposed a dynamic extension for the double Poisson model on continuous time, while Owen, (2011) proposed a discrete time random walk approach for both offensive and defensive parameters. In addition, Koopman and Lit, (2015) further extended the bivariate Poisson model into a state-space framework, allowing team abilities to vary according to a state vector.
A fundamentally different modelling approach has emerged due to the availability of large volumes of data, which has led to the development of a range of machine learning result-based techniques. These include artificial neural networks (ANNs), multivariate adaptive regression splines (MARS) (Friedman,, 1991), and ensemble learning methods such as random forests (Breiman,, 2001). Notably, the predictive performance of several random forests configurations has been examined and evaluated in the context of international football matches (Schauberger and Groll,, 2018; Groll et al.,, 2019, 2021, among others).
While the aforementioned methodologies provide robust frameworks for modelling football match outcomes, their predictive performance may be improved by incorporating additional historical information. Typically, one approach to potentially improve predictive accuracy is to integrate established rankings, such as the FIFA ranking, as additional model covariates. Specifically, these covariates are based on the quantitative measures of team strength from which the overall ranking is then derived. Notably, the algorithm underlying the FIFA ranking system has undergone a significant revision since 2018. The new algorithm takes into account not only the outcome of the single matches but also the strength level of teams before each match. For further details see Szczecinski and Roatis, (2022). The resulting algorithm offers benefits in terms of both simplicity and relative transparency. However, an alternative and particularly interesting methodology for obtaining a ranking, based on pairwise comparisons between teams, is the Bradley-Terry model (Bradley and Terry,, 1952). The model assigns a strength parameter to each team, and the odds of winning a match are determined by the ratio of these parameters. The estimated strengths can be employed to construct a rating system that reflects the ranking of teams based on the outcome of matches and the competitive interactions between teams.
In order to encompass a range of limitations, a number of extensions and generalizations to the original Bradley-Terry model have been developed over the years. For instance, Rao and Kupper, (1967) and Davidson, (1970) extended the model applicability to scenarios in which a draw is a possible outcome by including a novel parameter which affects the probability of a tie in a match. Springall, (1973) proposed a generalization of the Bradley-Terry model with team-dependent linear predictors. Davidson and Solomon, (1973) proposed a Bayesian version of the Bradley-Terry model, using a family of conjugate prior distributions to compute the posterior distribution of the log-strength parameters. Moreover, Leonard, (1977) suggested a more flexible Bayesian approach using non-conjugated multivariate Gaussian prior distributions for the log-strengths. Since these early works, there have been numerous contributions that have further extended the Bayesian framework for paired comparison models (Chen and Smith,, 1984; Caron and Doucet,, 2012; Whelan,, 2017; Osei and Davidov,, 2022; Wainer,, 2023, among others). The Bradley-Terry model can be further extended to account for situations where the order of comparisons can influence the outcome. A classic example is the home-field advantage in football, where the team playing at home may have a psychological or logistical advantage compared to the visiting team. Specifically, Beaver and Gokhale, (1975) and Davidson and Beaver, (1977) introduced additive and multiplicative order effects, respectively. Additionally, several extensions have been developed to model the dynamic variation of strengths over time. (Fahrmeir and Tutz,, 1994; Glickman,, 1999, 2001; Cattelan et al.,, 2012; Tian et al.,, 2023, among others).
Our work focuses on extending the well-known goal-based and result-based protocols by introducing alternative ranking measures for international football matches that could serve as a valuable computational routine for practitioners. The rest of the paper is organized as follows. Section 2 presents the theoretical framework, introducing the standard Bradley-Terry model and its Davidson extension for handling draws, followed by a discussion of the Bayesian approach. Furthermore, Section 3 describes the statistical goal-based methods and the machine learning result-based algorithms used in this study. In Section 4, we evaluate the application of these methodologies on the data from both the last World Cup and Africa Cup of Nations. Finally, Section 5 provides concluding remarks, outlining the limitations, advantages, and potential future research directions.
@@ 2 The Bradley-Terry Model
The Bradley-Terry model (Bradley and Terry,, 1952) is one of the most popular modelling techniques in a pairwise comparison context for ranking players or teams. The model assumes that each team $T_{k}$, with $k=1,\dots,N_{T}$, is characterized by a latent parameter, $\alpha_{k}>0$, representing its intrinsic strength, The outcome of any given comparison is modelled as an independent Bernoulli random variable, where the probability of each outcome is a function of the strengths of the teams involved. Specifically, for a match between team $T_{i}$ and team $T_{j}$, with $i\neq j=1,\dots,N_{T}$ , the probability that $T_{i}$ defeats $T_{j}$ in the n-th match, with $n=1,\dots,N$, is
@@
where $\alpha_{i}$ and $\alpha_{j}$ are the strength parameters of the teams involved in the match. These parameters are invariant to a multiplicative constant. Therefore, parameter identifiability is obtained by imposing a constraint such as $\sum_{k=1}^{N_{T}}\alpha_{k}=1$. Furthermore, the final ranking of the teams can be determined by sorting their respective strength parameters $\alpha_{k}$.
The model in (1) is commonly reparameterized by the logarithm of the strength parameters
@@ 2.1 Dealing with draws
$p_{ij}^{D}$ $=\dfrac{\exp(\gamma+(\psi_{i}+\psi_{j})/2)}{\exp(\psi_{i})+\exp(\psi_{j})+\exp(\gamma+(\psi_{i}+\psi_{j})/2)},$
$p_{ij}^{L}$ $=\dfrac{\exp(\psi_{j})}{\exp(\psi_{i})+\exp(\psi_{j})+\exp(\gamma+(\psi_{i}+\psi_{j})/2)}.$
Since the three-way process events are mutually exclusive, the following constraint is imposed $p_{ij}^{W}+p_{ij}^{D}+p_{ij}^{L}=1.$ It is worth noticing that if the draw parameter $\gamma$ increases towards +$\infty$+ ∞ then the probability of a tie $p_{ij}^{D}$ approaches one. Conversely, if $\gamma$ decreases towards $\infty$- ∞ then $p_{ij}^{D}$ approaches to zero. Finally, if $\gamma$ is equal to zero then $p_{ij}^{W}$, $p_{ij}^{D}$, and $p_{ij}^{L}$ depend solely on the strengths of the competing teams.
For the remainder of the paper, we will focus specifically on Davidson’s proposal for dealing with draws.
@@ 3 Statistical models and machine learning algorithms
This section describes the statistical models and machine learning algorithms employed to predict the outcomes of the considered competitions. Through a detailed analysis of goal-based models and result-based machine learning algorithms, this section aims to provide a comprehensive overview of the methodologies employed in the prediction of football matches, emphasizing their statistical foundations and practical implementations in sports analytics.
@@ 3.1 Goal-based models
Goal-based models assume that the number of goals scored in a match by each team follows a discrete distribution, typically two independent Poisson or a bivariate Poisson accounting for positive correlation. Thus, for each match, we need to consider the pair of counts $(X_{in},Y_{jn})$, for $i\neq j=1,\dots,N_{T}$ and $n=1,\dots,N$. The first count $X_{in}$ denotes the non-negative number of goals scored by the home team $T_{i}$ and the second count $Y_{jn}$ denotes the number of goals scored by the visiting team $T_{j}$, both in the n-th match. A simple double Poisson model is
@@
where $\lambda_{1n}$ and $\lambda_{2n}$ describe the expected number of goals for the home team and the away team, respectively. In particular, $\theta$ denotes a common baseline parameter, the parameters att and def represent the unknown attack and defense abilities for the home team $h_{n}$ and the away team $a_{n}$ in the n-th match. Furthermore, $\omega_{n}=(rank\_points_{h_{n}}-rank\_points_{a_{n}})$ captures the difference in FIFA ranking points (BTD relative log-strengths) between the home and away teams in the n-th match. Finally, the parameter $\phi$ tries to correct for the ranking points difference occurring between two competing teams. A sum-to-zero constraint (Baio and Blangiardo,, 2010) is imposed on the attack and defense parameters to ensure model identifiability.
@@
Poisson goal-based models may suffer from an underestimation of the number of draws, represented by the outcomes in the diagonal of the probability table. To address this issue, Karlis and Ntzoufras, (2009) introduced a zero-inflated model for favoring the draw outcome. The diagonal-inflated bivariate Poisson model is defined as follows
@@
where $\mu_{\operatorname{att}}$ and $\mu_{\operatorname{def}}$ are the mean for the initial attack and defense abilities, and $\sigma_{\operatorname{att}}$ and $\sigma_{\operatorname{def}}$ are their corresponding variances. As with the static models, the dynamic extension also imposes a sum-to-zero constraint on these random effects within each season for identifiability.
@@ 3.2 Result-based algorithms
Artificial neural networks (ANNs) represent a class of complex computational models inspired by the interconnected structure of neurons in the human brain. These models excel at processing information and learning from data through a layered architecture. Each layer comprises interconnected nodes that apply weights and biases to process inputs. The learning process involves adjusting these weights and biases to optimize the network’s performance. Specifically, ANNs proved successful in predicting football match outcomes by considering historical data that include a wide range of information, such as team performance metrics, match results, and even individual player statistics (Huang and Chang,, 2010; Hucaljuk and Rakipovic,, 2011; Danisik et al.,, 2018, among others).
@@
3.3 Computational procedure
In the statistical models and machine learning algorithms described in Sections 3.1 and 3.2 an additional predictor, determined by the difference in FIFA ranking points (BTD relative log-strengths) between the home team and the away team, is incorporated. Specifically, for the BTD relative log-strengths, this process involves initially fitting the Bayesian BTD model as described in Equation (4). Subsequently, the posterior median for each team’s log-strength parameter is computed. Finally, the difference in the posterior medians of the competing teams is included as an additional predictor in both the goal-based models and result-based algorithms.
The computational steps for integrating the Bayesian BTD relative log-strengths into the statistical models and machine learning algorithms are summarized in Algorithm 1.
Algorithm 1 Bayesian BTD computational steps
1:Fit the Bayesian BTD model as described in (4).
2:For $k=1,\ldots,N_{T}$ compute the posterior median for each team’s log-strength parameter $\psi_{k}$.
3:For $n=1,\ldots,N$ incorporate the difference in the posterior median of the competing teams’ log-strengths, $\omega_{n}=\psi_{h_{n}}-\psi_{a_{n}}$, as an additional predictor in both the goal-based models and result-based algorithms for the n-th match of the competition.